OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches 8 where f of x equals x plus 10 when x is less than 8 and f of x equals 10 minus x when x is greater than or equal to 8

OpenStudy (anonymous):

x+10 x less than 8 10-x x greater than or equal to 8

ganeshie8 (ganeshie8):

hey you should start using latex

OpenStudy (anonymous):

i have a few questions like this and really want to know how to solve problems like these

OpenStudy (anonymous):

latex?

ganeshie8 (ganeshie8):

\[f(x) = \left\{\begin{array}{} x+10&:& x \lt 8 \\10-x&:&x\ge 8\end{array}\right.\] like this ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

you want to find \(\lim\limits_{x\to 8} f(x)\)

OpenStudy (anonymous):

but it has lim with x-->8 under it

OpenStudy (anonymous):

You have to check if f is continuous at 8 and -8

ganeshie8 (ganeshie8):

we say the limit exists if you converge to the same value from both left and right sides of 8

ganeshie8 (ganeshie8):

plugin x=8 into both the given pieces of function the limit exists if you get same value

OpenStudy (anonymous):

so 18 x less than 8 and 2 x greater than or equal to 8

OpenStudy (anonymous):

they are both false

OpenStudy (anonymous):

Well they have to equal otherwise the limit does not exist

OpenStudy (anonymous):

by equalling each other you mean like getting 18 and 18 when I plug in the 8

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

oh okay thankssss aton !

OpenStudy (anonymous):

so for Find the indicated limit, if it exists. limit of f of x as x approaches 0 where f of x equals 5 x minus 9 when x is less than 0 and the absolute value of the quantity 2 minus x when x is greater than or equal to 0 the limit would not exist right?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i have gotten that for all my answers in my work

ganeshie8 (ganeshie8):

yes limit does not exist because the left and right limits converge to different values. graph makes it easy to see http://gyazo.com/e148f280b022d5186dc7258c1e017731

OpenStudy (anonymous):

Oh okay cause out of my 5 hw questions I have gotten so

OpenStudy (anonymous):

thanks again!

ganeshie8 (ganeshie8):

Notice as you approach 8 from left, the function goes to 18 but when you approach 8 from right, the function goes to 2 they dont converge to same value so we say \(\lim\limits_{x\to 8} f(x) = \text{Does Not Exist}\)